OpenStudy (anonymous):

MEDAL In triangle ABC, if segment BC is ne inch longer than segment AB, segment AC is 10 inches shorter than the sum of segments BC and AB, and the perimeter is 72 inches, find the length of segment BC. HALP MEH PLZ

OpenStudy (anonymous):

I meant one inch sorry

OpenStudy (anonymous):

length of segment BC (x): x + x - 1 + x + x - 1 - 10 = 72 4x = 84 x = 21 answer: 21 inches is the length of segment BC. proof (perimeter is 72 inches): = (21 - 1) + 21 + ([{21 - 1} + 21] - 10) = 20 + 21 + ([20 + 21] - 10) = 41 + (41 - 10) = 41 + 31 = 72

OpenStudy (anonymous):

thanks